OpenStudy (anonymous):

The function f(x)=ln(6−x) is represented as a power series. Find the first few coefficients in the power series.

OpenStudy (anonymous):

Yea because I have been stuck on this for a while. And could use any help.

OpenStudy (loser66):

to me, I consider the term (6-x) \[\frac{ 1 }{ 6(1- \frac{ x }{ 6 }) }= \frac{ 1 }{ 6 }\sum_{0}^{\infty} 1 + (\frac{ x }{ 6 })+(\frac{ x }{ 6 })^2+ (\frac{ x }{ 6 })^3+.......\] then take integral both sides I let 1/6 in the front aside, I calculate the next part. integral LHS = ln (6-x) and integral RHS you can take int few of them and stop at wherever you like

OpenStudy (loser66):

that's what I think, or you can multiple 1/6 into each term after the sum notation, and then take integral.

OpenStudy (anonymous):

Thank you that should help.

OpenStudy (loser66):

@sportymort let hear other's opinion

OpenStudy (loser66):

@Spacelimbus

OpenStudy (loser66):

he's off line, how ???

OpenStudy (anonymous):

Depends on what kind of topic this is based on, like usually this is an application of substitution. If we consider a Maclaurin Series then: \[\ln(1+x)= \sum_1^\infty \frac{ (-1)^\left( n+1 \right)x^n }{ n }\] Usually such kind of problems are faced first, nevertheless you can go from a substutiton from here. \[\ln(1+5-x)=\ln(6-x)=\sum_1^\infty \frac{ (-1)^\left(n+1 \right)(5-x)^n }{ n }\]

OpenStudy (loser66):

yep, much more neatly than mine.

OpenStudy (loser66):

anyway, he 's gone. I just want to study more from his problem.

OpenStudy (anonymous):

Yours was great @Loser66, that's the way how you would approach it manually, just make sure that you compute each successful derivative right. The results should remain the same though.